OpenStudy (anonymous):

Find the value of A^2-B^2 if A=2x^2+2x^-2 B=2x^2-2x^-2

OpenStudy (anonymous):

it's \((A-B)(A+B)\)

OpenStudy (anonymous):

\(A^2-B^2=(2x^2+2^{-2}-2x^2+2x^{-2})(2x^2+2^{-2}+2x^2-2x^{-2})=4(2^{-2})(2x^{2})\)

OpenStudy (anonymous):

if i'm not wrong.

OpenStudy (mathmale):

@AntarAzri : Our job is to guide others through problem solutions so that they're able to find their own answers. Please do not do others' work for them and provide free answers. You could give hints (such as "try squaring the formula for A).

OpenStudy (anonymous):

@mathmale Thank you i'm living a math flood :).

OpenStudy (anonymous):

find A+B and A-B then find the product.